OpenStudy (anonymous):

how do I convert 205 degrees 24' 55" to a decimal version of degrees?

OpenStudy (anonymous):

1degree =60' 1'=1/60 degree 24'=24/60degree this is the idea can be extended

OpenStudy (anonymous):

Ans is 205.41 degrees

OpenStudy (asnaseer):

as NotSObright indicated: 1 Degree = 60 Minutes = 60' 1 Minute = 60 Seconds = 60'' So, I Degree = 3600 Seconds Use these facts to convert the Minutes and Seconds to Degrees.

OpenStudy (y2o2):

I've an easy way: 1) divide the seconds by 60 (55/60) = 0.916666 2) add the minutes to the product , 24+0.91666 = 24.91666 3)divide it by 60 , 24.91666/60 = 0.415277 4) add the degrees to it , 205+0.415277 = 205.415277 5) That's all :D

OpenStudy (mertsj):

\[205^{o}24'55"=205^{o}+\frac{24}{60}^{o}+\frac{55}{3600}^{o}\]=

OpenStudy (mertsj):

\[205+.4+.015=205.415^{o}\]

OpenStudy (anonymous):

thank you guys!

OpenStudy (asnaseer):

yw